OpenStudy (anonymous):

Which equations in point-slope form are equations of the line that passes through the point (9, −2) and (−3, 6)? Choose exactly two answers that are correct. a. y+2=-2/3(x-9 b. y-9=-3/2(x+2) c. y+1=-3/2(x-4) d. y-6=-2/3(x+3)

OpenStudy (danjs):

do you recall the general form for point-slope y - y1 = m*(x - x1) where m= slope of the line (x1,y1) = the known point